13. A Quality Analyst is conduct a Failure Mode Effect Analysis on an Engine. The FMEA measures reveal an overall average of 705 cc. The average range is 6cc. If you use a sample size of 10, what are the upper and lower control limits for the mean and range? Upper Control Limit Lower Control Limit:
Answer #1

Overall average, X̅̅ = 705 cc

Average range, R̅ = 6 cc

Corresponding to sample size, n = 10, the control chart constants are:

A2 = 0.308

D3 = 0.223

D4 = 1.777

Upper Control Limit:

Mean: UCL = X̅̅ + A2*R̅ = 705 + 0.308*6 = 706.85

Range: UCLR = D4*R̅ = 1.777*6 = 10.66

Lower Control Limit:

Mean: LCL = X̅̅ - A2*R̅ = 705 - 0.308*6 = 703.15

Range: LCLR = D3*R̅ = 0.223*6 = 1.34


answered by: ANURANJAN SARSAM